A random sample of 24 wolf litters in Ontario, Canada, gave an average of x1 = 4.4 wolf pups per litter, with estimated sample standard deviation s1 = 1.0. Another random sample of 11 wolf litters in Finland gave an average of x2 = 2.8 wolf pups per litter, with sample standard deviation s2 = 1.3.

(a) Categorize the problem below according to parameter being estimated, proportion p, mean μ, difference of means μ1 – μ2, or difference of proportions p1 – p2. Then solve the problem.

μ

p

μ1 – μ2

p1 – p2

(b) Find an 90% confidence interval for μ1 – μ2, the difference in population mean litter size between Ontario and Finland. (Use 1 decimal place.)

lower limit

upper limit

(c) Examine the confidence interval and explain what it means in the context of this problem. Does the interval consist of numbers that are all positive? all negative? of different signs? At the 90% level of confidence, does it appear that the average litter size of wolf pups in Ontario is greater than the average litter size in Finland?

Because the interval contains only positive numbers, we can say that the average litter size of wolf pups is greater in Ontario.

Because the interval contains both positive and negative numbers, we can not say that the average litter size of wolf pups is greater in Ontario.

We can not make any conclusions using this confidence interval.

Because the interval contains only negative numbers, we can say that the average litter size of wolf pups is greater in Finland.
Answer #1

Ans n = 24, n2 = H, X12 4.4, A1 = 1.0 Tz = 2.0, 82= { 1,3 col Estimation parometers Horlo = 27-Yg = 4-4-2.0 (b) 90%CI for. 41